You may purchase 100 shares of Mun Tee ltd on a 55 percent margin when the shares are selling at K20 each. The Lusaka stock exchange broke charges you 10 percent annual interest,and commission are 3 percent of total stock value on both the purchase and the sale. If a year later you receive a K0.50 per share dividend and sell the stock for K27. What's your rate of return on investment?
1
Expert's answer
2021-09-17T16:16:25-0400

Number of shares = 100 shares

Margin percentage = 55%


Initial investment amount "= 100 \space shares \times 20 \times 55\% = 1100"

Interest expense on 45% of the investment from own funds

"= 20 ×

 (1−55 \%) × 100 × 10\% = 90"

Commission amount "= ( 20 + 27) \times 100 \times 3\% = 141"

Dividend amount "= 100 \times 0.50 = 50"

Amount received from sale "= 100 \times 27 = 2700"

Cost of shares "= 100 \times 20 = 2000"

Rate of return "=\frac{ 2700−2000+50−90−141}{

1100}"

Rate of return "= \frac{519}{

1100}


=0.4718 \space or\space 47.18\%"